# StructFlowBench: A Structured Flow Benchmark for Multi-turn Instruction Following

### v1

@@ 4.2 Main Results
Table 2 presents a comprehensive evaluation of 13 representative LLMs on StructFlowBench, covering four key metrics as well as assessments of structural constraints. The detailed results, categorized by intra-turn constraints and task types, are provided in the Appendix C.
@@ 4.4 Human Verification
We extracted 30 dialogues from the output of Qwen2-7B-Instruct and invited domain experts to conduct a comprehensive and detailed evaluation of the results. The experts rated the outputs using a binary scoring system. The results showed that the Kappa coefficient between GPT-4o’s evaluations and those of the experts was approximately 0.75. This indicates that utilizing advanced LLMs, like GPT-4o, to assess the quality of outputs from other models is a reliable approach, effectively reducing both subjective bias and the time costs associated with relying solely on human evaluation.
@@ 5 Conclusion
Currently, the structural flow in StructFlowBench is designed with a single linear relationship to facilitate analysis and data generation. For instance, if the third turn dialogue serves as both a recall structure to the first turn and a follow-up structure to the second turn, the current approach retains only the recall relationship while disregarding other structural dependencies. This simplification may limit the comprehensive modeling of hierarchical dialogue structures. Future work should extend the structural flow framework to simultaneously capture multiple coexisting dialogue relationships, thereby providing a more holistic representation of multi-turn dialogue complexity.
@@ Appendix A Details of Topics and Tasks
Appendix B Details of Constraints
The distribution of all constraints is detailed in Table 4, with the definitions of intra-turn constraints as follows:
Content Constraint: The response must strictly focus on the specified content scope and avoid any deviation from the topic.
Keyword/Element Constraint: The response must include specific words or elements as required.
Style Constraint: The response must be generated in a specific writing style, such as formal, humorous, poetic, etc.
Basic Format Constraint: The output must adhere to a specified basic format, such as JSON, XML, CSV, Table, Markdown, etc.
Quantity Format Constraint: The response must meet a precise requirement for the number of characters, words, sentences, or paragraphs as specified.
Template Format Constraint: The response must follow a predefined template structure, such as starting with a specific phrase, ending with a certain statement, or using a custom template provided by the user.
Situation Constraint: The response must be tailored to a given scenario or perspective, such as responding from a specific identity or context.
Inverse Constraint: The response must deliberately exclude or avoid certain constraints, such as not containing a specific keyword, not involving a particular element, or not using a certain language style.
@@ Appendix B Details of Constraints
Appendix C Detailed Results Categorized by Intra-turn Constraints and Task Types
@@
Appendix D Details of Prompts
@@
Appendix E Case of Data
Table 7 presents a sample case from StructFlowBench.
Structure "source": "c1","target": "c2","relation": "follow-up"
"source": "c1","target": "c3","relation": "recall"
"source": "c3","target": "c4","relation": "unrelatedness"
"source": "c4","target": "c5","relation": "refinement"
Summarized Prompts "c1" : "The user asks the assistant, role-playing as a financial advisor, to provide a general strategy for a young professional who wants to start saving for retirement."
"c5": "The user modify the detail level in last round’s prompt to request a deeper dive into the unique instruments used in each genre for better understanding of their sounds."
Complete Dialogue "name": "c1",
"name": "c5",
Check Lists "name":"c1"
"name":"c5"
"Basic Format Constraint":"Is the response formatted as a table?"
"Refinement Constraint":"Is the c5 conversation a refinement of c4 conversation?"
Table 7: An example of synthetic data.
Appendix F Details of Models
All the details about the evaluated models are provided in Table 8.
Model Model Link
GPT GPT-4o https://platform.openai.com/docs/models#gpt-4o
Claude Claude-3.5-Sonnet https://docs.anthropic.com/en/docs/about-claude/models
Gemini Gemini-1.5-Pro https://ai.google.dev/gemini-api/docs/models/gemini?hl=en#gemini-1.5-pro
Deepseek DeepSeek-v3 https://huggingface.co/deepseek-ai/DeepSeek-V3
DeepSeek-R1-Distill-Qwen-7B https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B
DeepSeek-R1-Distill-Llama-8B https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B
Qwen Qwen2.5-14B-Instruct https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
Qwen2.5-7B-Instruct https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
GLM GLM-4-9B-Chat https://huggingface.co/THUDM/glm-4-9b-chat
Yi Yi-6B-Chat https://huggingface.co/01-ai/Yi-6B-Chat
LLAMA Llama-3.1-8B-Instruct https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
Mistral Mistral-7B-Instruct-v0.3 https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3
Phi Phi-3.5-mini-instruct https://huggingface.co/microsoft/Phi-3.5-mini-instruct
Table 8: Model Links.

### v2

@@ 4.2 Main Results
Table 2 presents a comprehensive evaluation of 13 representative LLMs on StructFlowBench, covering four key metrics as well as assessments of structural constraints. The detailed results, categorized by intra-turn constraints and task types, are provided in the Appendix B.
@@ 4.3.1 Complex Scenario Suitability Study
4.3.2 Fail Case Analysis of Refinement
Given that none of the models achieved optimal performance on the Refinement structure, we conducted an in-depth analysis of failure cases specific to this category. Our analysis suggests that the high error rate in Refinement tasks stems from the model’s difficulty in distinguishing refinement from follow-up instructions, particularly in multi-turn settings. In refinement scenarios, only specific constraints should be updated, while others—often introduced in earlier turns—must remain intact. However, models frequently overlook these earlier constraints, failing to preserve them in the revised responses. This forgetting behavior leads to violations of the intended dialogue structure and undermines the overall satisfaction of constraints. A concrete example of such a failure is illustrated in Figure 5.
Figure 5: A Fail Case of Refinement.
@@ 4.4 Human Verification
We extracted 30 dialogues from the output of Qwen2.5-7B-Instruct and invited two domain experts to conduct a comprehensive and detailed evaluation of the results. The task requirements provided to the human evaluators are shown in Appendix Figure 11. Specifically, the experts were instructed to assess adherence to constraints through a binary evaluation protocol. Each constraint was assigned a “Yes” for full alignment and a “No” for any violation. The results showed that the Kappa coefficient between GPT-4o’s evaluations and those of the experts was approximately 0.75. This indicates that utilizing advanced LLMs, like GPT-4o, to assess the quality of outputs from other models is a reliable approach, effectively reducing both subjective bias and the time costs associated with relying solely on human evaluation.
@@ 5 Conclusion
The current StructFlowBench is more of an exploration of new directions in evaluating multi-turn dialogue instruction following, rather than a targeted effort to increase the difficulty of the evaluation set. As a result, leading models tend to achieve near-maximum scores on many metrics. Future work can enhance the difficulty by extending the number of dialogue turns, introducing more constraints per turn, and incorporating more challenging tasks such as complex reasoning and retrieval-augmented generation (RAG).
In addition to increasing task difficulty, another important direction lies in improving the structural design itself. Currently, the structural flow in StructFlowBench is designed with a single linear relationship to facilitate analysis and data generation. For instance, if the third turn dialogue serves as both a recall structure to the first turn and a follow-up structure to the second turn, the current approach retains only the recall relationship while disregarding other structural dependencies. This simplification may limit the comprehensive modeling of hierarchical dialogue structures. Future work should extend the structural flow framework to simultaneously capture multiple coexisting dialogue relationships, thereby providing a more holistic representation of multi-turn dialogue complexity.
@@
Acknowledgments
The authors would like to thank the anonymous reviewers for their valuable comments. This work is supported by the National Key Research and Development Program of China (No.2023YFF0905400), the National Natural Science Foundation of China (No.U2341229) and the Reform Commission Foundation of Jilin Province (No.2024C003).
@@ Appendix A Details of Topics and Tasks
Constraint Name Definition Example
Content Constraint The response must strictly focus on the specified content scope and avoid any deviation from the topic. Is the response focused on recommending a graphics card for a gaming PC?
Keyword/Element Constraint The response must include specific words or elements as required. Must contain the word “Artificial Intelligence” in your answer.
Style Constraint The response must be generated in a specific writing style, such as formal, humorous, poetic, etc. Please write a report in a formal style.
Basic Format Constraint The output must adhere to a specified basic format, such as JSON, XML, CSV, Table, Markdown, etc. Please output the following data in JSON format.
Quantity Format Constraint The response must meet a precise requirement for the number of characters, words, sentences, or paragraphs as specified. Please provide an answer in no more than 100 words.
Template Format Constraint The response must follow a predefined template structure, such as starting with a specific phrase, ending with a certain statement, or using a custom template provided by the user. Please follow the template below to generate your content.
Situation Constraint The response must be tailored to a given scenario or perspective, such as responding from a specific identity or context. Imagine you are an experienced doctor and respond to the following health-related questions.
Inverse Constraint The response must deliberately exclude or avoid certain constraints, such as not containing a specific keyword, not involving a particular element, or not using a certain language style. Make sure your response does not involve duscrimination or politics.
Follow-up Constraint The response must reflect a follow-up structure, building upon the previous turn by incorporating elements from the user’s or AI’s prior input to ensure dialogue continuity. Is the T2 turn a follow-up of T1?
Refinement Constraint The response must conform to a refinement structure, where the user modifies or clarifies their immediately preceding prompt to improve relevance or accuracy while maintaining context. Is the T3 turn a refinement of T2?
Expansion Constraint The response must align with an expansion structure, where a main theme is introduced and related subtopics are explored across multiple turns with thematic continuity. Is the [T2,T3] turns an expansion of the T1?
Summary Constraint The response must follow a summary structure, consolidating information from multiple previous turns into a coherent and concise overview with clarity and completeness. Is the T4 turn a summary of [T1,T2,T3]
Recall Constraint The response must adhere to a recall structure, referencing content from two or more turns earlier to re-establish context or seek clarification, ensuring long-range coherence. Is the T5 turn a recall of T2?
Table 4: Constraint System of StructFlowBench
@@
Appendix B Detailed Results Categorized by Intra-turn Constraints and Task Types
@@
Structure "source": "T1","target": "T2","relation": "follow-up"
"source": "T1","target": "T3","relation": "recall"
"source": "T3","target": "T4","relation": "unrelatedness"
"source": "T4","target": "T5","relation": "refinement"
Summarized Prompts "T1" : "The user asks the assistant, role-playing as a financial advisor, to provide a general strategy for a young professional who wants to start saving for retirement."
"T5": "The user modify the detail level in last round’s prompt to request a deeper dive into the unique instruments used in each genre for better understanding of their sounds."
Complete Dialogue "name": "T1",
"name": "T5",
Check Lists "name":"T1"
"name":"T5"
"Basic Format Constraint":"Is the response formatted as a table?"
"Refinement Constraint":"Is the T5 conversation a refinement of T4 conversation?"
Table 8: An example of synthetic data.
Model Paremeters (Billions) Model Link
GPT GPT-4o $\sim 200$ https://platform.openai.com/docs/models#gpt-4o
Claude Claude-3.5-Sonnet $\sim 175$ https://docs.anthropic.com/en/docs/about-claude/models
Gemini Gemini-1.5-Pro $\sim 175$ https://ai.google.dev/gemini-api/docs/models/gemini?hl=en#gemini-1.5-pro
Deepseek DeepSeek-v3 671 https://huggingface.co/deepseek-ai/DeepSeek-V3
DeepSeek-R1-Distill-Qwen-7B 7 https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B
DeepSeek-R1-Distill-Llama-8B 8 https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B
Qwen Qwen2.5-14B-Instruct 14 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
Qwen2.5-7B-Instruct 7 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
GLM GLM-4-9B-Chat 9 https://huggingface.co/THUDM/glm-4-9b-chat
Yi Yi-6B-Chat 6 https://huggingface.co/01-ai/Yi-6B-Chat
LLAMA Llama-3.1-8B-Instruct 8 https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
Mistral Mistral-7B-Instruct-v0.3 7 https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3
Phi Phi-3.5-mini-instruct 3.8 https://huggingface.co/microsoft/Phi-3.5-mini-instruct
Table 9: Overview of Selected Large Language Models with Parameter Sizes and Reference Links
Appendix C Details of Prompts
@@
Figure 11: Task Requirements For Human Evaluators